# Asymptotic behavior of eigenvalues of large rank perturbations of large random matrices

### v2

@@ Abstract
The paper is concerned with deformed Wigner random matrices. These matrices are closely connected with Deep Neural Networks (DNNs): weight matrices of trained DNNs could be represented in the form $R+S$, where $R$ is random and $S$ is highly correlated. The spectrum of such matrices plays a key role in rigorous underpinning of the novel pruning technique based on Random Matrix Theory. Mathematics has been done only for finite-rank matrix $S$. However, in practice rank may grow. In this paper we develop asymptotic analysis for the case of growing rank.
@@ 1 Introduction
Random Matrix Theory (RMT) is a classical theory that has been developing for more than 70 years. Initially, RMT arose from problems in nuclear physics and found its applications in mathematics, physics, finance, and many other disciplines. Recently, new problems have been arising from the area of Machine Learning. Indeed, often the weight matrices of Deep Neural Networks (DNNs) are initialized randomly. Moreover, modern DNNs have large weight matrices, which is why their spectral properties can be described by asymptotic behavior of $N\times N$ random matrices as $N$ goes to infinity.
@@ 1.1 Formulation of the problem
where $R$ is a rectangular matrix whose entries are i.i.d. random variables with zero mean and variance $\sigma^{2}$ and $S$ is a real symmetric non-random (or random, but strongly correlated) matrix. The present paper is concerned with similar but simpler $N\times N$ deformed Gaussian Orthogonal Ensemble matrices of the same form
@@
where $R$ is restricted to be a quadratic real symmetric matrix whose entries are Gaussian random variables $\mathcal{N}(0,\sigma^{2})$ and $S$ is also a real symmetric non-random matrix.
@@
$\mu_{N}(\Delta)=\frac{1}{N}\sum_{j=1}^{N}\mathbbm{1}_{\Delta}(\lambda_{j})$ (1.3)
where $\Delta$ is an interval in $\mathbb{R}$, $\lambda_{j}$ – eigenvalues of $X$ and $\mathbbm{1}_{\Delta}$ is a characteristic function of the set $\Delta$. For a probability measure $\tau$ on $\mathbb{R}$, denote by $g_{\tau}$ its Stieltjes transform defined for $z\in\mathbb{C}\setminus\operatorname{supp}(\tau)$ by
@@
In 1972 Pastur established that the ESD $\mu$ of the matrices (1.2) converges as $N\to\infty$, assuming the ESD $\nu$ of $S$ converges to the measure $\nu_{0}$ [5], see also [1]. It was also shown that the Stieltjes transform $g_{\mu_{0}}(z)$ of the limiting ESD $\mu_{0}$ of the matrix $W$ defined (1.2) satisfies the equation
@@
In the particular case when $\nu_{0}$ is a delta-measure at zero, $\mu_{0}$ coincides with a well-known semicircular law $\mu_{sc}$ [8]
@@
Moreover, if rank of $S$ is finite, i.e. does not depend on $N$, Péché established the behavior of the largest eigenvalue $\lambda_{1}(W)$ in [7]. It was shown that if $\lim\limits_{N\to\infty}\lambda_{1}(S)=\theta_{1}$ then almost surely
$\lim\limits_{N\to\infty}\lambda_{1}(W)=\begin{cases}2\sigma,\quad\text{if }\theta_{1}<\sigma,\\ \Phi(\theta_{1})=\theta_{1}+\frac{\sigma^{2}}{\theta_{1}},\quad\text{if }\theta_{1}\geq\sigma,\end{cases}$ (1.8)
where
@@
Let $\mu$ and $\nu$ be the ESDs of $W$ and $S$ respectively. We impose the following assumptions on the matrix $S$ and its spectrum.
@@
The crucial distinction from the previous results is Assumption 2. All the previously obtained theorems assume that the deterministic matrix $S$ has a finite rank $r(N)=r^{\prime}$ independent of $N$.
In [3] Berlyand et al provided mathematical underpinning to the MP-pruning algorithm in the following setting: the weight matrix to be pruned can be represented as a sum of random and low-rank (‘‘signal’’) components with non-intersecting spectra (1.1). The asymptotical behavior of the spectrum of the matrices (1.1) plays a key role in bounding the difference between weight matrices before and after pruning, although the mathematics is done only for finite rank. However, restriction $\lim_{N\to\infty}r(N)<\infty$ severe bounds applicability of pruning. In practice it seems that a growing rank $r(N)\to\infty$, as $N\to\infty$, is in better agreement with numerically computed weight matrices of DNNs. Numerical simulations show that the number of singular values of weight matrices may grow, see Subsection 1.2 and Figure 1. This paper develops asymptotical analysis for the case of growing rank.
@@ 1.2 Motivation: weight matrices of DNNs
Let us consider a DNN for solving a classification problem. The setup of a classification problem is as following. There is a set of some objects $\mathcal{S}$. Each object $s\in\mathcal{S}$ belongs to one of $K$ classes. Given the object, one wants to determine its class. We look for a solution in the form of a function $\varphi\colon\mathcal{S}\to[0;1]^{K}$, interpreting components of $\varphi(s)$ as probabilities for $s$ to belong to each class.
@@
where each $f_{k}:\mathbb{R}^{N_{k-1}}\to\mathbb{R}^{N_{k}}$, $f_{k}(x,\alpha^{k})=W_{k}x+\beta_{k}$ is an affine function, $W_{k}$ is a $N_{k}\times N_{k-1}$ weight matrix, $\beta_{k}\in\mathbb{R}^{N_{k}}$ is a bias vector, $\lambda:\mathbb{R}\to\mathbb{R}$ is a nonlinear activation function, which is applied component-wise (for example, absolute value), $\sigma$ is a function which normalizes vector’s components to probabilities. The entries of $W_{k}$ and the biases $\beta_{k}$ are the parameters $\alpha^{k}$ of a DNN. The parameters are optimized by minimizing the loss function, using, for example, stochastic gradient descent. See [2] for more details.
@@
Pruning parameters of a DNN is the process of removing certain weights (set them equal to zero) from the network in a way which preserves the overall performance. The goal is to eliminate components that contribute little to the output of the DNN function, resulting in a more sparse and more efficient architecture. Pruning parameters of deep neural networks is an important task because it reduces complexity of the function, leading to smaller, faster, and more efficient models. This not only decreases memory and computational requirements — crucial for deployment on resource-constrained devices — but can also improve generalization by preventing overfitting. In [3] Berlyand et al provided mathematical underpinning to the MP-pruning algorithm in the following setting: the weight matrix to be pruned can be represented as a sum of random and low-rank (‘‘signal’’) components with non-intersecting spectra (1.1). The asymptotical behavior of the spectrum of the matrices (1.1) plays a key role, although the mathematics is done only for finite rank.
@@
The fixed low-rank structure of the ‘‘signal’’ matrix allows for certain analytical tools to be employed, leading to provable guarantees on network behavior, particularly the bounds on accuracy and loss after pruning. However, numerical experiments reveal that this low-rank assumption does not always hold in practice. Realistic weight matrices often have a spectrum that has a ‘‘signal’’ part with the number of non-zero eigenvalues increasing with respect to the matrix size (see Figure 1). See Appendix A for more details. These observations suggest that the existing theoretical results, while insightful, are not fully representative of the behavior seen in real-world networks.
@@
Figure 1: Numeric simulation of a DNN with 3 layers. Show the dependence of the number of non-zero eigenvalues of the ‘‘signal’’ matrix, so called spikes, on the size of the matrix.
@@
In this work, we study a mathematical problem applicable in a simplified yet insightful setting: a DNN with two consecutive layers of equal size. This structure implies that the weight matrix connecting these layers is square. Furthermore, we impose an additional assumption of symmetry on the weight matrix, which not only aligns with certain model architectures but also simplifies the theoretical analysis.
@@
Motivated by the gap between existing rigorous proofs and numerical results, our goal is to generalize the existing theoretical framework to accommodate symmetric weight matrices without a fixed-rank constraint. By relaxing the low-rank assumption, we aim to provide a more comprehensive understanding of the network’s accuracy and the behavior of the loss function. This generalization not only brings the theory closer to practical scenarios but also raises interesting mathematical questions about the behavior of random matrices beyond the low-rank-perturbation setting. Through this work, we seek to bridge the divide between theory and numerical computations, contributing to a more robust foundation for analyzing deep learning algorithms.
@@ 2 Main results
In this section we present our main results. We first consider the measure $\frac{N}{r}(\mu-\mu_{0})$, which may be interpreted as an ESD of eigenvalues outside the bulk. The following theorem proves existence of the limiting measure and describes it.
@@
Let $W$ be a random matrix of the form (1.2). Let $\mu$ and $\nu$ be the ESDs of matrices $W$ and $S$, respectively, $\mu_{0}$ be the limit of $\mu$ as $N\to\infty$, and let Assumptions 1–3 hold. Then a signed measure $\frac{N}{r}(\mu-\mu_{0})$ converges weakly in distribution, as $N\to\infty$, to a non-random measure $\mu_{1}$. Moreover, if $\mu_{1}$ has a density, then for any measurable set $\Delta$ which does not intersect with $\operatorname{supp}\mu_{0}$
@@
The next theorem describes the asymptotic behavior of individual eigenvalues. To introduce the theorem, we need more notation. For any matrix $X$ let its eigenvalues be
@@
Let $W$ be a random matrix of the form (1.2). Let $\mu$ and $\nu$ be the ESDs of matrices $W$ and $S$ respectively and let Assumptions 1–3 hold. Additionally, we assume $\operatorname{supp}\nu_{1}^{+}=[a,b]$, $\operatorname{dist}(a,\operatorname{supp}\nu_{0})>0$, $\lambda_{j}(S)$, $1\leq j\leq r$, are outside $\operatorname{supp}\nu_{0}$ and
$\lim_{N\to\infty}\max_{1\leq j\leq N}\operatorname{dist}(\lambda_{j}(S),\operatorname{supp}\nu_{0}\cap[a,b])=0;$ (2.3)
$\lim_{N\to\infty}\max_{1\leq j\leq r}\operatorname{dist}(\lambda_{j}(S),[a,b])=0.$ (2.4)
@@
[…]
@@
    Theorem 2.2 agrees well with the result (1.8) obtained by Péché [7].
The paper is organized as follows. Section 3 is concerned with a pre-limiting equation for the Stieltjes transform of the ESD (like (1.5), but before taking the limit $N\to\infty$). We improve the error term in the equation for using it further in the proof. In Sections 4 and 5 we present proofs of Theorems 2.1 and 2.2 respectively.
@@ 3 Equation for the Stieltjes transform
Let $W$ be defined in (1.2) and let the ESD $\nu$ of $S$ converge weakly in distribution to $\nu_{0}$ as $N\to\infty$. Then the Stieltjes transform $g_{\mu}(z)$ satisfies the following equation
@@
$UWU^{*}=\frac{1}{\sqrt{N}}URU^{*}+\Sigma$
has the same spectrum as $W$, and $URU^{*}$ has the same distribution as $R$. Thus, everywhere below in this proof we treat $S$ as a diagonal matrix.
@@
Taking inverse of a block matrix, one obtains
@@
and $W^{(i)}$ is a $(N-1)\times(N-1)$ matrix obtained from $W$ by removing its $i$th column and its $i$th row, $\bm{w}_{i}$ is the $i$th column of $W$ without $W_{ii}$.
Since $S$ is diagonal, we have $\bm{w}_{i}=\frac{1}{\sqrt{N}}\bm{r}_{i}$. Let us transform the last term in (3.5)
@@
The proof is given in Appendix.
@@
Proof of Lemma 3.2.
Let $\operatorname{\mathbf{E}}_{i}$ be an expectation w.r.t. $\Re R_{ij}$ and $\Im R_{ij}$, $j=1,\dotsc,N$. We have
$\operatorname{\mathbf{E}}_{i}\left\{r_{1}\overline{r_{1}}\right\}=\dfrac{1}{N^{2}}\sum_{j_{1}\neq k_{1};\ j_{1},k_{1}\neq i}^{N}\sum_{j_{2}\neq k_{2};\ j_{2},k_{2}\neq i}^{N}G_{j_{1}k_{1}}^{(i)}\overline{G}_{j_{2}k_{2}}^{(i)}\operatorname{\mathbf{E}}_{i}\left\{R_{ij_{1}}\overline{R}_{ik_{1}}R_{ij_{2}}\overline{R}_{ik_{2}}\right\}.$ (3.18)
$\operatorname{\mathbf{E}}_{i}\left\{R_{ij_{1}}\overline{R}_{ik_{1}}R_{ij_{2}}\overline{R}_{ik_{2}}\right\}=\operatorname{\mathbf{E}}_{i}\left\{R_{ij_{1}}\right\}\operatorname{\mathbf{E}}_{i}\left\{\overline{R}_{ik_{1}}R_{ij_{2}}\overline{R}_{ik_{2}}\right\}=0.$ (3.19)
$\begin{split}\operatorname{\mathbf{E}}_{i}\left\{r_{1}\overline{r_{1}}\right\}&=\dfrac{1}{N^{2}}\sum_{\begin{subarray}{c}j_{1}\neq k_{1}\\ j_{1},k_{1}\neq i\end{subarray}}G_{j_{1}k_{1}}^{(i)}\overline{G}_{k_{1}j_{1}}^{(i)}\operatorname{\mathbf{E}}_{i}\left\{\left\lvert R_{ij_{1}}R_{ij_{2}}\right\rvert^{2}\right\}\\ &\quad{}+\dfrac{1}{N^{2}}\sum_{\begin{subarray}{c}j_{1}\neq k_{1}\\ j_{1},k_{1}\neq i\end{subarray}}^{N}\left\lvert G_{j_{1}k_{1}}^{(i)}\right\rvert^{2}\operatorname{\mathbf{E}}_{i}\left\{R_{ij_{1}}^{2}\overline{R}_{ik_{1}}^{2}\right\}\\ &\leq\frac{2}{N^{2}}\operatorname{Tr}G^{(i)}(G^{(i)})^{*}\leq\frac{2}{N}\left\lVert G^{(i)}\right\rVert^{2}\leq\frac{2}{N\left\lvert\Im z\right\rvert^{2}}.\end{split}$ (3.20)
$\begin{split}\operatorname{\mathbf{E}}_{i}\left\{r_{2}\overline{r_{2}}\right\}&=\dfrac{1}{N^{2}}\sum_{j_{1}\neq i}\sum_{j_{2}\neq i}G_{j_{1}j_{1}}^{(i)}\overline{G}_{j_{2}j_{2}}^{(i)}\operatorname{\mathbf{E}}_{i}\left\{(\left\lvert R_{ij_{1}}\right\rvert^{2}-\sigma^{2})(\left\lvert R_{ij_{2}}\right\rvert^{2}-\sigma^{2})\right\}\\ &=\dfrac{1}{N^{2}}\sum_{j\neq i}\left\lvert G_{jj}^{(i)}\right\rvert^{2}\operatorname{\mathbf{E}}_{i}\left\{(\left\lvert R_{ij}\right\rvert^{2}-\sigma^{2})^{2}\right\}\leq\frac{c_{4}}{N^{2}}\operatorname{Tr}G^{(i)}(G^{(i)})^{*}\leq\frac{c_{4}}{n\left\lvert\Im z\right\rvert^{2}},\end{split}$ (3.21)
$\operatorname{\mathbf{E}}\left\{r_{3}\overline{r_{3}}\right\}=\operatorname{Var}\left\{\frac{1}{N}\operatorname{Tr}G^{(i)}\right\}\leq\frac{C}{N},$ (3.22)
$\left\lvert\operatorname{Tr}G^{(i)}-\operatorname{Tr}G\right\rvert\leq\frac{1}{\left\lvert\Im z\right\rvert}$, (3.23)
one has $r_{4}\leq\frac{C}{N}$. ∎

### v3

@@ Abstract
The paper is concerned with deformed Wigner random matrices. These matrices are closely related to Deep Neural Networks (DNNs): weight matrices of trained DNNs could be represented in the form $R+S$, where $R$ is random and $S$ is highly correlated. The spectrum of such matrices plays a key role in rigorous underpinning of the novel pruning technique based on Random Matrix Theory. In practice, the spectrum of the matrix $S$ can be rather complicated. In this paper, we develop an asymptotic analysis for the case of full rank $S$ with increasing number of outlier eigenvalues.
@@ 1 Introduction
Random Matrix Theory (RMT) is a classical theory that has been developing for more than 70 years. Initially, RMT arose from problems in nuclear physics and found applications in mathematics, physics, finance, and many other disciplines. Recently, new problems have been arising in the area of Machine Learning. Indeed, often the weight matrices of Deep Neural Networks (DNNs) are initialized randomly. Moreover, modern DNNs have large weight matrices, which is why their spectral properties can be described by the asymptotic behavior of $N\times N$ random matrices as $N$ goes to infinity.
@@ 1.1 Formulation of the problem
where $R$ is a rectangular matrix whose entries are i.i.d. random variables with zero mean and variance $\sigma^{2}$ and $S$ is a real non-random matrix (or random, but strongly correlated).
The present paper is concerned with a similar but simpler $N\times N$ deformed Wigner Ensemble matrices of the form
@@
where $R$ is restricted to be a square real symmetric matrix whose entries are i.i.d. random variables with zero mean and variance
$\operatorname{\mathbf{E}}\left\{R_{jk}^{2}\right\}=(1+\delta_{jk})\sigma^{2},$ (1.3)
$\operatorname{\mathbf{E}}$ denotes the expectation and $\delta_{jk}$ is $1$ if $j=k$ and 0 otherwise. $S$ is a real symmetric non-random matrix.
@@
$\mu_{N}(\Delta)=\frac{1}{N}\sum_{j=1}^{N}\mathbbm{1}_{\Delta}(\lambda_{j}(X))$ (1.4)
where $\Delta$ is an interval in $\mathbb{R}$,
$\lambda_{1}(X)\geq\lambda_{2}(X)\geq\dotsb\geq\lambda_{N}(X)$ (1.5)
are eigenvalues of $X$ and $\mathbbm{1}_{\Delta}$ is a characteristic function of the set $\Delta$. For a probability measure $\tau$ on $\mathbb{R}$, denote by $g_{\tau}$ its Stieltjes transform defined for $z\in\mathbb{C}\setminus\operatorname{supp}(\tau)$ by
@@
In 1972 Pastur established that the ESD $\mu$ of the matrices (1.2) converges as $N\to\infty$, assuming that the ESD $\nu$ of $S$ converges to the measure $\nu_{0}$ [7], see also [1]. It was also shown that the Stieltjes transform $g_{\mu_{0}}(z)$ of the limiting ESD $\mu_{0}$ of the matrix $W$ defined in (1.2) satisfies the equation
@@
In the particular case of $\nu_{0}$ being a delta-measure at zero, $\mu_{0}$ coincides with a well-known semicircular law $\mu_{sc}$ [10]
@@
The individual asymptotic behavior of the eigenvalues $\lambda_{j}(W)$ depends on the individual asymptotic behavior of the eigenvalues $\lambda_{j}(S)$. The eigenvalues of $S$ may either converge to the bulk (i.e., $\operatorname{supp}\nu_{0}$) or not. Let us assume that the largest $r(N)$ eigenvalues $\lambda_{j}(S)$ do not converge to the bulk, but that the other do.
$\max_{1\leq j\leq r(N)}\operatorname{dist}(\lambda_{j}(S),\operatorname{supp}\nu_{0})>\varepsilon>0;$ (1.10)
$\lim_{N\to\infty}\max_{r(N)<j\leq N}\operatorname{dist}(\lambda_{j}(S),\operatorname{supp}\nu_{0})=0.$ (1.11)
The eigenvalues satisfying the property (1.10) are called spikes or outliers.
For $R$ being a matrix from a Gaussian Unitary Ensemble and for a low rank matrix $S$ ($\lambda_{j}(S)=0$ for $j>r(N)$) with $\lim\limits_{N\to\infty}\frac{r(N)}{N}=0$, Péché established the behavior of the largest eigenvalue $\lambda_{1}(W)$ in [9]. In particular, for fixed $r(N)=r^{\prime}$ it was shown that if $\lim\limits_{N\to\infty}\lambda_{1}(S)=\theta_{1}$ then almost surely
$\lim\limits_{N\to\infty}\lambda_{1}(W)=\begin{cases}2\sigma,\quad\text{if }\theta_{1}<\sigma,\\ \theta_{1}+\frac{\sigma^{2}}{\theta_{1}},\quad\text{if }\theta_{1}\geq\sigma.\end{cases}$ (1.12)
The general case where $R$ is drawn from the Wigner Ensemble was considered in [3] and [4]. In order to describe these results, let us introduce a function
@@
Note that if $\nu_{0}$ is a delta-function at zero, then $\Phi(\theta_{1})$ coincides with the second line in (1.12). Capitaine et al [3] dealt with fixed $r(N)=r^{\prime}$ and with $S$ of general form. They showed the following. Let $\lim\limits_{N\to\infty}\lambda_{j}(S)=\theta_{j}$, $j=1,\dotsc,r^{\prime}$ and let the common distribution of the entries of $R$ be symmetric and satisfy a Poincaré inequality. Then $\lambda_{j}(W)$ converges almost surely to $\Phi(\theta_{j})$ if $\Phi^{\prime}(\theta_{j})>0$, and to the bulk otherwise. In [4] Huang showed a similar convergence of eigenvalues, but in the case where $r(N)\to\infty$, $r(N)\ll N$, and $S$ is a diagonal low rank matrix, namely, all the eigenvalues of $S$ except the spikes are zeros.
Let $\mu$ and $\nu$ be the ESDs of $W$ and $S$, respectively, and let $S$ have exactly $r=r(N)$ non-zero eigenvalues. Denote
$\eta_{N}(\Delta)=\frac{1}{r}\sum_{\lambda_{j}(S)\neq 0}\mathbbm{1}_{\Delta}(\lambda_{j}(S)).$ (1.14)
We impose the following assumptions on the matrix $S$ and its spectrum.
@@
The crucial distinction from the previous results is that we simultaneously consider the general form of $\nu_{0}$ and $r(N)\to\infty$. In [3] the measure $\nu_{0}$ is of general form, but the deterministic matrix $S$ has a finite rank $r(N)=r^{\prime}$ independent of $N$, while in [4] $r(N)$ goes to infinity but $\nu_{0}$ is a delta-measure at zero.
@@ 1.2 Motivation: weight matrices of DNNs
Let us consider a DNN to solve a classification problem. The setup of a classification problem is as follows. There is a set of some objects $\mathcal{S}$. Each object $s\in\mathcal{S}$ belongs to one of $K$ classes. Given the object, one wants to determine its class. We search for a solution in the form of a function $\varphi\colon\mathcal{S}\to[0;1]^{K}$, interpreting the components of $\varphi(s)$ as probabilities that $s$ belongs to each class.
@@
where each $f_{k}:\mathbb{R}^{N_{k-1}}\to\mathbb{R}^{N_{k}}$, $f_{k}(x,\alpha^{k})=W_{k}x+\beta_{k}$ is an affine function, $W_{k}$ is an $N_{k}\times N_{k-1}$ weight matrix, $\beta_{k}\in\mathbb{R}^{N_{k}}$ is a bias vector, $\lambda:\mathbb{R}\to\mathbb{R}$ is a nonlinear activation function, which is applied component-wise (for example, absolute value), $\sigma$ is a function which normalizes vector’s components to probabilities. The entries of $W_{k}$ and the biases $\beta_{k}$ are the parameters $\alpha^{k}$ of a DNN. The parameters are optimized by minimizing the loss function, using, for example, stochastic gradient descent. See [2] for more details.
@@
Pruning parameters of a DNN is the process of removing certain weights (set them equal to zero) from the network in a way which preserves the overall performance. The goal is to eliminate components that contribute little to the output of the DNN function, resulting in a more sparse and more efficient architecture. Pruning parameters of deep neural networks is an important task because it reduces the complexity of the function, leading to smaller, faster, and more efficient models. This not only decreases memory and computational requirements — crucial for deployment on resource-constrained devices — but can also improve generalization by preventing overfitting. In [5] Berlyand et al. provided mathematical underpinning to the pruning algorithm in the following setting: the weight matrix to be pruned can be represented as a sum of random and low-rank (‘‘signal’’) components with non-intersecting spectra (1.1). They call it Marchenko-Pastur(MP)- pruning as the key step is setting the singular values of the weight matrix smaller than the right edge of the support of Marchenko-Pastur distribution to zero, which helps remove unnecessary randomness from the DNN. The asymptotical behavior of the spectrum of matrices (1.1) plays a key role.
@@
The fixed low-rank structure of the ‘‘signal’’ matrix allows certain analytical tools to be employed, leading to provable guaranties on network behavior, particularly the bounds on accuracy and loss after pruning. However, numerical experiments reveal that this low-rank assumption does not always hold in practice. Realistic weight matrices often have a spectrum that has a ‘‘signal’’ part with the number of non-zero eigenvalues increasing with respect to the matrix size (see Figure 1). See Appendix A for more details. These observations suggest that the existing theoretical results, while insightful, are not fully representative of the behavior seen in real-world networks.
@@
Figure 1: Numeric simulation of a DNN with 3 layers. Show the dependence of the number of outlier eigenvalues of the ‘‘signal’’ matrix, so called spikes, on the size of the matrix.
@@
In this work, we study a mathematical problem that is applicable in a simplified yet insightful setting: a DNN with two consecutive layers of equal size. This structure implies that the weight matrix connecting these layers is square. Furthermore, we impose an additional assumption of symmetry on the weight matrix, which is not applicable for DNNs directly, but simplifies the theoretical analysis.
@@
Motivated by the gap between existing rigorous proofs and numerical results, our goal is to generalize the existing theoretical framework to accommodate symmetric weight matrices without a low-rank constraint or a fixed number of spikes. By relaxing these assumptions, we aim to provide a more comprehensive understanding of the accuracy of the network and the behavior of the loss function. This generalization brings the theory closer to practical scenarios, namely the ‘‘bulk decay’’ case (see [6, Figure 12(b) and 12(d)]), where the distribution of eigenvalues looks not like the MP-distribution plus some spikes, but like the dilated MP distribution plus some spikes. Moreover, this generalization raises interesting mathematical questions about the behavior of random matrices beyond the low-rank perturbation setting. Through this work, we seek to bridge the divide between theory and numerical computations, contributing to a more robust foundation for analyzing deep learning algorithms.
@@ 2 Main results
In this section, we present our main results. We first consider the measure $\frac{N}{r}(\mu-\mu_{0})$, which can be interpreted as an ESD of eigenvalues outside the bulk. The following theorem proves the existence of the limiting measure and describes it.
@@
Let $W$ be a random matrix of the form (1.2). Let the third and fourth moments of the entries of $R$ be finite. Let $\mu$ and $\nu$ be the ESDs of the matrices $W$ and $S$, respectively, $\mu_{0}$ be the limit of $\mu$ as $N\to\infty$, and let Assumptions 1–3 hold. Then a signed measure $\frac{N}{r}(\mu-\mu_{0})$ converges weakly in distribution, as $N\to\infty$, to a non-random measure $\mu_{1}$. Moreover, if $\mu_{1}$ has a density, then for any measurable set $\Delta$ that does not intersect with $\operatorname{supp}\mu_{0}$
@@
The next theorem describes the asymptotic behavior of individual eigenvalues. To introduce the theorem, we need more notation. For any matrix $X$, let its eigenvalues be
@@
Let $W$ be a random matrix of the form (1.2). Let the third and fourth moments of the entries of $R$ be finite. Let $\mu$ and $\nu$ be the ESDs of matrices $W$ and $S$, respectively, and let Assumptions 1–3 hold. Additionally, we assume that $\operatorname{supp}\nu_{1}^{+}=[\omega_{\mu_{0}}(a),\omega_{\mu_{0}}(b)]$, $\operatorname{dist}(\omega_{\mu_{0}}(a),\operatorname{supp}\nu_{0})>0$, $\lambda_{j}(S)$, $1\leq j\leq r$, are outside $\operatorname{supp}\nu_{0}$ and
$\lim_{N\to\infty}\max_{1\leq j\leq N}\operatorname{dist}(\lambda_{j}(S),\operatorname{supp}\nu_{0}\cup[\omega_{\mu_{0}}(a),\omega_{\mu_{0}}(b)])=0;$ (2.3)
$\lim_{N\to\infty}\max_{1\leq j\leq r}\operatorname{dist}(\lambda_{j}(S),[\omega_{\mu_{0}}(a),\omega_{\mu_{0}}(b)])=0.$ (2.4)
@@
    If the sequence $j(N)$ is chosen such that $\lim_{N\to\infty}\lambda_{j(N)}(S)=\theta$, then $\lambda_{j(N)}(W)\to\Phi(\theta)$, as $N\to\infty$.
@@
    Theorem 2.2 agrees well with the results obtained in [3, 4, 9].
The paper is organized as follows. Section 3 is concerned with a pre-limiting equation for the Stieltjes transform of the ESD (like (1.7), but before taking the limit $N\to\infty$). We improve the error term in the equation to use it further in the proof. In Sections 4 and 5 we present proofs of Theorems 2.1 and 2.2, respectively.
@@ 3 Equation for the Stieltjes transform
In this section, we prove that an error term in the pre-limiting equation for the Stieltjes transform of the ESD has the order $O(N^{-1})$. We do it in two steps: first, for the case of deformed Gaussian Orthogonal Ensemble (GOE)
$\tilde{W}=N^{-1/2}H+S,$ (3.1)
where $N^{-1/2}H$ is drawn from GOE, and then in the general case. Proposition 3.1 covers the first step.
@@
Let $\tilde{W}$ be defined in (3.1). Let the ESD $\nu$ of $S$ converge weakly to $\nu_{0}$ as $N\to\infty$. Then the Stieltjes transform $g_{\mu}(z)$ satisfies the following equation
@@
$U\tilde{W}U^{*}=\frac{1}{\sqrt{N}}UHU^{*}+\Sigma$
has the same spectrum as $\tilde{W}$, and $UHU^{*}$ has the same distribution as $H$. Thus, it is enough to prove the claim for the matrices $W$ of the form
$W=N^{-1/2}H+\Sigma.$ (3.3)
In a standard way, we write the Stieltjes transform in terms of the resolvent of $W$
@@
Taking the inverse of a block matrix, one obtains
@@
and $W^{(i)}$ is an $(N-1)\times(N-1)$ matrix obtained from $W$ by removing its $i$th column and its $i$th row, $\bm{w}_{i}$ is the $i$th column of $W$ without $W_{ii}$.
Since $\Sigma$ is diagonal, we have $\bm{w}_{i}=\frac{1}{\sqrt{N}}\bm{r}_{i}$. Let us transform the last term in (3.7)
@@
The proof of the lemma is given in Appendix B.
@@
Proposition 3.3.
Let $W$ be defined in (1.2). Let $R$ be a square real symmetric matrix whose entries are i.i.d. random variables with zero mean and variance $\sigma^{2}$, and let the ESD $\nu$ of $S$ converge weakly to $\nu_{0}$ as $N\to\infty$. Then the Stieltjes transform $g_{\mu}(z)$ satisfies the following equation
Proof.
Introduce a parametrized family of matrices $W(t)$ in the following way:
$W(t)=\dfrac{1}{\sqrt{N}}R(t)+S$ (3.21)
$R(t)=\sqrt{t}R+\sqrt{1-t}H.$ (3.22)
where entries of $R$ are i.i.d. random variables, with mean zero and variance $\sigma^{2}$, and $H$ is a Gaussian matrix (GOE), whose entries have mean zero and variance $\sigma^{2}$.
Note that for $t=0$, $R(0)=H$ with Gaussian entries, and we proved (3.20) for $t=0$, in Proposition 3.1. Here, we need to show that the same equation holds for $W(1)$, when the random matrix $R(1)=R$ is a Wigner matrix.
Introduce $G(t,z)$, the resolvent of $W(t)$
$G(t,z)=(W(t)-z)^{-1}$ (3.23)
and the Stieltjes transform $g_{\mu}(t,z)$
$g_{\mu}(t,z)=\dfrac{1}{N}\operatorname{\mathbf{E}}\{\operatorname{Tr}G(t,z)\}=\dfrac{1}{N}\sum\operatorname{\mathbf{E}}\left\{G_{ii}(t,z)\right\}.$ (3.24)
We prove the following lemma, first
Lemma 3.4.
$\left|\dfrac{dg_{\mu}(t,z)}{dt}\right|=O\left(\dfrac{1}{N}\right)$ uniformly in $t\in[0,1]$.
The proof of the lemma is given in Appendix B. Now, Lemma 3.4 implies
$\int_{0}^{1}\dfrac{dg_{\mu}(t,z)}{dt}dt=g_{\mu}(1,z)-g_{\mu}(0,z)=O\left(\dfrac{1}{N}\right)$ (3.25)
Thus,
$g_{\mu}(1,z)=g_{\mu}(0,z)+O\left(\dfrac{1}{N}\right)$ (3.26)
where $g_{\mu}(1,z)=g_{\mu}(z)$ is an object of interest in (3.20) and for $g_{\mu}(0,z)$ the same statement was proven in Prop.3.1. As the (3.2) and (3.20) both hold up to an order $O\left(\dfrac{1}{N}\right)$, and $g_{\mu}(1,z)$ and $g_{\mu}(0,z)$ differ from each other by the term of the same order, we conclude that (3.20) holds for $g_{\mu}(1,z)$. ∎